Hey! Can someone PLEASE help me?!? I honestly have no clue what i'm doing and i'm so worried that i'm gonna fail♥...Here is the problem: Solve each equation if x and y are whole numbers. 2xy+5=15 and xy=4-x.
Help would be greatly appreciated!♥
♥Jenna♥

Answer by ikdeep(226)

2xy+5=15 ...............(1)
xy=4-x...................(2)
its very simple..by substitution method,just substitute the value of xy from (1) into (2)..and you get............

2(4-x)+5 = 15
now solve brackets..and we get....
8 - 2x + 5 = 15
now put together like variables.....
8+5 - 2x = 15
13 - 2x = 15
or -2x = 15 - 13
-2x = 2
x = 2/-2
x = 1/-1
and to put into standarad form we get....
x = -1
now to get the value of y you can substitute the value of x = -1 in either (1)or(2)....but for easy calculation ,,here i took (2)..and we get...
xy=4-x...................(2)
-1y = 4-(-1)
-1y = 4+1
-1y = 5
y = 5/-1
or to put into standarad we get,,,,
y = -5
,therefore the value of x = -1 and y = -5

to verify your answer,,,simply put the values of x and y either in (1)or (2) or in both ,,,you will always get LHS= RHS....


please feel free to revert back for any further queries.